The solubility of benzoic acid at 0 degrees celcius is 1.7 g/L and at 90 degrees celcius is 45.5g/L. how much water do you need to dissolve 0.91g of benzoic acid at 90 degrees Celcius? How much of the original benzoic acid will remian in the solution once the temperature is reduced to 0 degrees celcius?
1
Expert's answer
2012-05-31T08:48:16-0400
The solubility of benzoic acid at 90 degrees celcius is 45.5 g/L. We need
0.91/45.5 = 0.02 L (20 mL) of water.
At 0 degrees celcius the solution
contain 0.91/1.7=0.54 g of benzoic acid. Thus, 0.91-0.54=0.37g of benzoic acid
remains in the solution at 0 degrees celcius.